### v3

# Optimistic Multi-Agent Policy Gradient

Abstract
Relative overgeneralization (RO) occurs in cooperative multi-agent learning tasks when agents converge towards a suboptimal joint policy due to overfitting to suboptimal behaviors of other agents. No methods have been proposed for addressing RO in multi-agent policy gradient (MAPG) methods although these methods produce state-of-the-art results. To address this gap, we propose a general, yet simple, framework to enable optimistic updates in MAPG methods that alleviate the RO problem. Our approach involves clipping the advantage to eliminate negative values, thereby facilitating optimistic updates in MAPG. The optimism prevents individual agents from quickly converging to a local optimum. Additionally, we provide a formal analysis to show that the proposed method retains optimality at a fixed point. In extensive evaluations on a diverse set of tasks including the Multi-agent MuJoCo and Overcooked benchmarks, our method outperforms strong baselines on 13 out of 19 tested tasks and matches the performance on the rest.
1 Introduction
Multi-agent reinforcement learning (MARL) is a promising approach for many cooperative multi-agent decision making applications, such as those found in robotics and wireless networking (Busoniu et al., 2008). However, despite recent success on increasingly complex tasks (Vinyals et al., 2019; Yu et al., 2022), these methods can still fail on simple two-player matrix games as shown in Figure 1. The underlying problem is that in cooperative tasks, agents may converge to a suboptimal joint policy when updating their individual policies based on data generated by other agents’ policies which have not converged yet. This phenomenon is called relative overgeneralization (Wiegand, 2004) (RO) and has been widely studied in tabular matrix games (Claus & Boutilier, 1998; Lauer & Riedmiller, 2000; Panait et al., 2006) but remains an open problem in state-of-the-art MARL methods (De Witt et al., 2020; Yu et al., 2022; Kuba et al., 2022).
The cause for RO can be understood intuitively. In a cooperative multi-agent system, the common reward derives from the joint actions. From the perspective of a single agent, an optimal individual action may still incur low joint reward due to non-cooperative behaviors of other agents. This is common in cooperative tasks when iteratively optimizing individual policies, especially at the beginning of training since individual agents have not learned to act properly to cooperate with others. It is particularly challenging in tasks with a large penalty for incorrect joint actions, such as the pitfalls in the climbing matrix game in Figure 1 leading often to agents that prefer a suboptimal joint policy.
Existing techniques to overcome the pathology share the idea of applying optimistic updating in Q-learning with different strategies to control the degree of optimism (Lauer & Riedmiller, 2000) such as in hysteretic Q-learning (Matignon et al., 2007) or in lenient agents (Panait et al., 2006). However, these methods are designed based on Q-learning and only tested on matrix games (Matignon et al., 2012) with tabular Q representation. Although optimism successfully helps overcome RO problem and converges to global optima in tabular tasks, it could amplify the overestimation problem when combined with DQN (Van Hasselt et al., 2016) with function approximation (Omidshafiei et al., 2017; Palmer et al., 2018; Rashid et al., 2020a), as verified in our experiments. On the other hand, recent multi-agent policy gradient (MAPG) methods have achieved state-of-the-art performance on popular MARL benchmarks (De Witt et al., 2020; Yu et al., 2022; Sun et al., 2023; Wang et al., 2023) but still suffer from the RO problem, converging to a low value local optimum in simple matrix games. The above drawbacks motivate this work to investigate whether optimism can be applied to MAPG methods and whether it can boost performance further by mitigating the RO problem.
Our contribution is threefold: (1) To our knowledge, we are the first to investigate the application of optimism in MAPG methods. We propose a general, yet simple, framework to incorporate optimism into the policy gradient computation in MAPG. Specifically, we propose to clip the advantage values when updating the policy. For completeness, we extend the framework to include a hyperparameter to control the degree of optimism, resulting in a Leaky ReLU function (Maas et al., 2013) to reshape the advantage values. (2) We provide a form analysis to show that the proposed method retains optimality at a fixed point. (3) In our experiments111Source Code: https://github.com/wenshuaizhao/optimappo, the proposed OptiMAPPO algorithm successfully learns global optima in matrix games and outperforms both recent state-of-the-art MAPG methods MAPPO (Yu et al., 2022), HAPPO, HATRPO (Kuba et al., 2022), and existing optimistic methods in complex domains.
2 Related work
In this section, we discuss classic optimistic methods and optimistic DQN based approaches. We also discuss recent MAPG methods that yield state-of-the-art performance on common benchmarks. Optimistic Thompson sampling (O-TS) (Hu et al., 2023) is also discussed as it utilizes a similar clipping technique to improve exploration for stochastic bandits. General multi-agent exploration methods are related and introduced. For completeness, we further discuss advantage shaping in single-agent settings.
Classic Optimistic Methods
To the best of our knowledge, distributed Q-learning (Lauer & Riedmiller, 2000) proposes the first optimistic updating in independent Q-learning, where -values are only updated when the value increases. Lauer & Riedmiller (2000) also provides brief proof that the proposed optimism-based method converges to the global optimum in deterministic environments. To handle stochastic settings, hysteretic Q-learning (Matignon et al., 2007) adjusts the degree of optimism by setting different learning rates for the Q values of actions with different rewards. The frequency maximum Q value heuristic (FMQ) (Kapetanakis & Kudenko, 2002) considers changing the action selection strategies during exploration instead of modifying the updating of Q values. Lenient agent (Panait et al., 2006; Wei & Luke, 2016) employs more heuristics to finetune the degree of optimism by initially adopting an optimistic disposition and gradually transforming into average reward learners.
Optimistic Deep Q-Learning
Recently, several works have extended optimistic methods and lenient agents to Deep Q-learning (DQN). Dec-HDRQN (Omidshafiei et al., 2017) and lenient-DQN (Palmer et al., 2018) apply hysteretic Q-learning and leniency to DQN, respectively. Optimistic methods have also been combined with value decomposition methods. Weighted QMIX (Rashid et al., 2020a) uses a higher weight to update the Q-values of joint actions with high rewards and a lower weight to update the values of suboptimal actions. FACMAC (Peng et al., 2021) improves MADDPG (Lowe et al., 2017) by taking actions from other agents’ newest policies while computing the Q values, which can also be regarded as optimistic updating as it is natural to suppose the newer policies of other agents would generate better joint Q-values. Even though the optimism has been applied to DQN-based methods, in common benchmarks they are usually outperformed by recent MAPG methods (Yu et al., 2022). This unsatisfying performance of optimistic DQN methods can be attributed to the side effect of overestimation of Q-values, which we also empirically verify in our experiments.
Multi-Agent Policy Gradient Methods
COMA (Foerster et al., 2018) is an early MAPG method in parallel with value decomposition methods (Sunehag et al., 2017; Rashid et al., 2020b). It learns a centralized on-policy function and utilizes it to compute the individual advantage for each agent to update the policy. However, until IPPO and MAPPO (De Witt et al., 2020; Yu et al., 2022) which directly apply single-agent PPO (Schulman et al., 2017) into multi-agent learning, MAPG methods show significant success on popular benchmarks. The strong performance of these methods might be credited to the property that individual trust region constraint in IPPO/MAPPO can still lead to a centralized trust region as in single-agent PPO (Sun et al., 2023). Nonetheless, HAPPO/HATRPO (Kuba et al., 2022) and A2PO (Wang et al., 2023) further improve the monotonic improvement bound by enforcing joint and individual trust region, while with the cost of sequentially updating each agent. However, even with a strong performance on popular benchmarks, these methods don’t explicitly consider the relative overgeneralization problem during multi-agent learning and can still converge to a suboptimal joint policy.
Multi-Agent Exploration Methods
Since RO can be seen as a special case of the general exploration problem we discuss below general MARL exploration methods. Similar to single-agent exploration, Rashid et al. (2020b) and Hu et al. (2021) use noise for exploration in multi-agent learning. However, as shown by (Mahajan et al., 2019) noise-based exploration can result in suboptimal policies. For coordinated exploration, multi-agent variational exploration (Mahajan et al., 2019) conditions the joint Q-value on a latent state. Jaques et al. (2019) maximizes the mutual information between agent behaviors for coordination. This may nevertheless still lead to a sub-optimal joint strategy (Li et al., 2022). Zhao et al. (2023) propose conditionally optimistic exploration (COE) which augments agents’ Q-values by an optimistic bonus based on a global state-action visitation count of preceding agents. However, COE is designed for discrete states and actions and is difficult to scale up to complex tasks with continuous state and action spaces. Cooperative multi-agent exploration (CMAE) (Liu et al., 2021) only counts the visitations of states in a restricted state space to learn an additional exploration policy. However, the restricted space selection is hard to scale up and thus CMAE fails to show performance improvement on widely used MARL benchmarks. Our method aims to mitigate the specific RO problem with state-of-the-art MAPG methods in order to further boost their performance on complex tasks.
Optimistic Thompson Sampling
Thompson sampling is a popular method for stochastic bandits (Russo et al., 2018). Conceptually, Thompson Sampling plays an action according to the posterior probability distribution of the optimal action. The key idea for optimistic Thompson sampling (O-TS) (Hu et al., 2023) is to clip the posterior distribution in an optimistic way to ensure that the sampled models are always better than the empirical models (Chapelle & Li, 2011). O-TS shares similar heuristics with our method, that is, there is no need to decrease a prediction in Thompson sampling, or no need to explicitly decrease the action probability in policy updates. However, it is not straightforward to apply O-TS to model-free policy gradient methods. Our method achieves optimism by a novel advantage clipping instead of the posterior distribution clipping in O-TS.
Advantage Shaping
A few works have explored advantage shaping in single-agent RL settings. PPO-CMA (Hämäläinen et al., 2020) tackles the prematurely shrinking variance problem in PPO by clipping or mirroring the negative advantages in order to increase exploration and eventually converges to a better policy. Self-imitation learning (SIL) (Oh et al., 2018) learns an off-policy gradient from replay buffer data with positive advantages in addition to the regular on-policy gradient. However, different from these existing works, we are motivated by solving the RO problem in multi-agent learning.
3 Background
We begin by introducing our problem formulation. Following this, we present the concept of optimistic Q-learning as it forms the basis of hysteresis and leniency based approaches which differ in their ways regulating the degree of optimism.
3.1 Problem Formulation
We mainly study the fully cooperative multi-agent sequential decision-making tasks which can be formulated as a multi-agent Markov decision process (MMDP) (Boutilier, 1996) consisting of a tuple , where is the set of agents. At time step of Dec-MDP, each agent observes the full state in the state space of the environment, performs an action in the action space from its policy . The joint policy consists of all the individual policies . The environment takes the joint action of all agents , changes its state following the dynamics function and generates a common reward for all the agents. is a reward discount factor. The agents learn their individual policies and maximize the expected return: . Since we train independent policy networks for each agent, we use and interchangeably as they all include the full state information in our experiments.
3.2 Optimistic Q-learning
We explain the idea of optimistic Q-learning (Matignon et al., 2007) based on hysteretic Q-learning. While regular Q-learning update assigns the same learning rate to both negative and positive updates, hysteretic Q-learning assigns a higher weight to the positive update of the Q value, i.e., when the right-hand side (RHS) of Equation 1 has a higher value than the left-hand side (LHS). In our experiments on the baseline hysteretic Q-learning, it is equivalent to only set the weight for negative updates, leaving the positive update with the default learning rate.
Optimism with tabular Q-learning has been demonstrated effective to solve the RO problem. However, with function approximation, the optimistic update could exacerbate the overestimation problem (Van Hasselt et al., 2016) of deep Q-learning and thus fail to improve the underlying methods, which is also shown in our experiments. To our knowledge, the application of optimism in MAPG methods has not been explored and it remains unclear how to facilitate optimism in policy gradient methods and how much improvement it can escort.
4 Method
Aware of the importance of optimism in solving RO problem and the limitation of optimistic Q-learning, we instead propose a principled way to apply optimism to the recent MAPG methods. The algorithm we use in experiments is instantiated based on MAPPO (Yu et al., 2022) but note that the proposed framework can be further applied to other advantage actor-critic (A2C) based MARL methods as shown in Appendix E.
4.1 Optimistic MAPPO
In MAPPO (Yu et al., 2022), each agent learns a centralized state value function , and the individual policy is updated via maximizing the following objective
where is the clipping threshold and is the importance ratio between the current policy and the previous policy used to generate the data,
The advantage is usually estimated by the generalized advantage estimator (GAE) (Schulman et al., 2016) defined as
and denotes the TD error
While PPO (Schulman et al., 2017) adopts the clipping operation to constrain the policy change in order to obtain the guaranteed monotonic improvement, the policy update is similar to common A2C (Mnih et al., 2016) methods. The policy is improved by increasing the actions with positive advantage values and decreasing the others with negative advantages. However, the actions currently with negative advantages might be the optimal action and the current negation comes from the currently suboptimal teammates, not from the suboptimality of the actions. In tasks with a severe relative overgeneralization problem, optimal actions are often not recovered by simple exploration strategies and the joint policy converges to a suboptimal solution.
In order to overcome the relative overgeneralization problem in MAPG methods, the proposed optimistic MAPPO (OptiMAPPO) applies a clipping operation to reshape the estimated advantages (Schulman et al., 2016). OptiMAPPO optimizes the agents’ policies by maximizing the following new objective
where denotes that negative advantage estimates are clipped to zero while positive advantage values remain unchanged. is the same clipping operation in PPO (Schulman et al., 2017). The proposed advantage clipping operation allows to be optimistic to temporarily suboptimal actions incurred by the RO problem and facilitates individual agents to converge to a better joint policy. The implementation can be as straightforward as a single-line modification of underlying MAPG methods. However, as demonstrated by our experiments, the effectiveness of optimism is significantly enhanced compared to optimistic Q-learning based methods.
Extension to Leaky ReLU operation
The proposed clipping operation can be seen as a special case of a Leaky ReLU (LR) operation of advantage values where there is a hyperparameter to control the degree of optimism, . Our clipping operation is the case when , while recovers the original MAPG methods. In our experiments, we find that the performance is improved more while setting lower , i.e. higher optimism. However, we argue that such an extension could be beneficial in stochastic reward environments as discussed in hysteretic Q-learning (Matignon et al., 2007) and lenient agents methods (Palmer et al., 2018; Matignon et al., 2012). The extension allows us to control the degree of optimism in a finer granularity. We leave the study in stochastic environments as future work and this paper is primarily focused on the first to investigate the effectiveness of optimism in MAPG methods.
4.2 Analysis
The following analysis shows formally from an operator view that the proposed algorithm retains optimality at a fixed point.
The operator view fits our method in the context of policy gradient methods. As shown in (Ghosh et al., 2020), the policy update in policy gradient methods can be seen as two successive operations from improvement operator and projection operator . Such an operator view connects both Q-learning and vanilla policy gradient by accounting to different , which is detailed in the Appendix A.
We show that the proposed advantage clipping forms a new improvement operator that retains optimality at a fixed point of the operators. To simplify the analysis, we only take the clipping operation instead of the extended Leaky ReLU operation, as it transforms the advantage values into non-negative values, satisfying the valid probability distribution requirement in the operator view derivation. Based on the clipped advantage, the new policy gradient is
Note that we use and interchangeably. It corresponds to the new improvement operator formulated as
where and are defined as:
The projection operator remains the same as in vanilla policy gradient. With the clipped advantage, the optimal policy is a fixed point of the operators as in the vanilla policy gradient. The property is shown in the following Proposition 4.1 and proven in the Appendix B.
Proposition 4.1.
is a fixed point of ,
where denotes the function composition notation. It means the latter operator is first evaluated and then its output will be used by the former operator .
5 Experiments
We compare our method with the following strong MAPG baselines on both illustrative matrix games and complex domains including Multi-agent MuJoCo and Overcooked:
MAPPO directly applies singe-agent PPO into multi-agent settings while with a centralized critic. Despite the lack of theoretical guarantee, MAPPO has achieved tremendous success in a variety of popular benchmarks.
HAPPO is the first-order emulation algorithm of HATRPO that follows the idea of PPO.
We further compare the proposed optimistic MAPG method with existing optimistic Q-learning based methods. In Multi-agent MuJoCo with continuous action space, we include the recent FACMAC (Peng et al., 2021) as our optimistic baseline which overcomes the RO problem by using the other agents’ newest policy to update individual policy. Hysteretic DQN (Matignon et al., 2007; Omidshafiei et al., 2017) is used as our optimistic baseline in the Overcooked domain which has discrete action space.
The RO problem can be seen as a special category of the general exploration problem. Therefore, we also investigate whether existing multi-agent exploration methods can solve the RO problem. NA-MAPPO (Hu et al., 2021) is a general exploration method by injecting noise into the advantage estimates. MAVEN (Mahajan et al., 2019) facilitates coordinated exploration by conditioning the joint Q-value on a latent state. CMAE (Liu et al., 2021) learns a separate exploration policy based on the count of the visitations of states.
Our method is implemented based on MAPPO and we use the same hyperparameters as MAPPO in all tasks. For HAPPO and HATRPO, we follow their original implementation and hypermeters but align the learning rate and the number of rollout threads to have a fair comparison. The implementation details including the pseudo-code of OptiMAPPO can be found in Appendix C.
5.1 Repeated Matrix Games
Even though with small state and action spaces, the climbing and penalty matrix games (Claus & Boutilier, 1998), as shown in Figure 1, are usually hard to obtain the optimal joint solution without explicitly overcoming the relative overgeneralization problem. The matrix games have two agents which select the column and row index of the matrix respectively. The goal is to select the correct row and column index to obtain the maximal element of a matrix. In the penalty game, is the penalty term and we evaluate for . The lower value of , the harder for agents to identify the optimal policy due to the growing risk of penalty. Following (Papoudakis et al., 2021), we set constant observation and the episode length of repeated games as .
[…]
| Climbing | ||||
| Penalty k=0 | ||||
| Penalty k=-25 | ||||
| Penalty k=-50 | ||||
| Penalty k=-75 | ||||
| Penalty k=-100 |
To better understand the RO problem and the role of being optimistic in these cooperative tasks, Figure 1 compares the learning process with and without the optimistic update on the Climbing task. Initially, both agents uniformly assign probability on each index. In each step , the agents update their individual policy distribution based on (Abdolmaleki et al., 2018; Peters et al., 2010). is calculated as , is fixed as 2 and is the payoff at row , column . From Figure 1, it is clear to see after the first update step, without handling the RO issue (the first row), the joint policy quickly assigns a high probability on a sub-optimal solution and assigns low probabilities on the rest, while optimistic update avoids it. This is important in cooperative tasks to prevent premature convergence. As we can see after 10 update steps, the baseline method converges to a sub-optimal solution by selecting the number while the optimistic update successfully finds the global solution.
In Table 1, we compare our method with other MAPG baselines. We can see that OptiMAPPO with optimistic update achieves the global optima, while the baseline without considering the relative overgeneralization problem converge to the local optima. The performance of more popular MARL methods on matrix games can be found in Appendix D.
5.2 Multi-agent MuJoCo (MA-MuJoCo)
In this section, we investigate whether the proposed OptiMAPPO can scale to more complex continuous tasks and how it compares with the state-of-the-art MAPG methods. MA-MuJoCo (Peng et al., 2021) contains a set of complex continuous control tasks which are controlled by multiple agents jointly. The evaluation results of the selected three tasks are presented in Figure 2 and the full results on all the tasks are left in the Appendix G. We observe that OptiMAPPO obtains clearly better asymptotic performance in most tasks compared to the baselines.
In the Humannoid Standup task which needs agents to coordinate well to stand up, the baselines experience drastic oscillation during learning while OptiMAPPO increases much more stably. We would like to attribute the oscillation in the baselines to the RO problem as the strong simultaneous coordination between agents is necessary to stand up. We also compare the maximum episodic returns during evaluation in Appendix, where our method outperforms the baselines by a large margin on most tasks.
We observe that at the beginning of the training, our method learns more slowly than the baselines on the MA-MuJoCo tasks, however, this is not as evident on the Overcooked tasks. This is because the advantage clipping disregards the information of the data with negative advantages, which can be seen as a cost to converge to a better solution. However, on Overcooked, since the task spaces are smaller compared to MA-MuJoCo tasks, the samples are sufficient in each update.
Comparison with FACMAC
We include FACMAC as our optimistic baseline (Peng et al., 2021) for the continuous action space tasks. FACMAC has been proposed to improve MADDPG (Lowe et al., 2017) by solving the RO problem with a centralized gradient estimator. Specifically, FACMAC samples all actions from agents’ current policies when evaluating the joint action-value function, which can be seen as one way to achieve optimism since the newest policy would generate higher Q estimation. As the results listed in Table 2, OptiMAPPO significantly outperforms FACMAC, which shows that our method can better employ the advantage of optimism and achieve stronger performance than the current optimistic method.
5.3 Overcooked
Overcooked (Carroll et al., 2019; Yu et al., 2023) is a fully observable two-player cooperative game that requires the agents to coordinate their task assignment to accomplish the recipe as soon as possible. We test OptiMAPPO on tasks with different layouts. To succeed in these games, players must coordinate to travel around the kitchen and alternate different tasks, such as collecting onions, depositing them into cooking pots, and collecting a plate.
The comparisons with both recent MAPG algorithms and existing optimism baseline are shown in Figure 3, where our method consistently achieves similar or better performance on all tasks. For example in the Random3 task, We observe that the baseline methods either consistently converge to a suboptimal policy (in HAPPO), or show a big variance between different training (in MAPPO). In the rendered videos, the suboptimal policy in HAPPO can only learn to assign one agent to deliver, while our method successfully learns to rotate in a circle to speed up the delivery. We speculate that the gap between our method and baselines is due to the presence of subtle RO problems in certain tasks. The baseline methods can only rely on naive exploration to find the optimal joint policy, which is susceptible to the RO problem. However, our method with advantage shaping can effectively overcome these issues and converge more stably. In Section 5.6, even though the optimistic baseline hysteretic DQN also utilizes optimism to overcome the RO problem, it can incur drastic overestimation and thus fails to improve the performance, as shown in our experiments.
5.4 Comparison with Exploration Methods
The results of MAVEN (Mahajan et al., 2019) and NA-MAPPO (Hu et al., 2021) on the matrix games are shown in Table 3, where both methods fail to solve the RO problem. The experiments show that general exploration methods following the principle of being optimistic in the face of uncertainty (Munos et al., 2014; Imagawa et al., 2019) may not be able to solve the RO problem. Our method works by being optimistic to the suboptimal joint actions instead of unseen states or actions. Comparisons with NA-MAPPO in the MA-MuJoCo domain can be found in Appendix F.
[…]
| Climbing | |||
| Penalty k=0 | |||
| Penalty k=-25 | |||
| Penalty k=-50 | |||
| Penalty k=-75 | |||
| Penalty k=-100 |
We also compare our algorithm with CMAE (Liu et al., 2021) on Penalized Push-Box. This benchmark modifies the original Push-Box task in (Liu et al., 2021) by injecting penalty to agents when the agents are not coordinated to push box at the same time, thereby presenting the RO issue. Table 4 shows that both our method and CMAE can successfully overcome the RO problem and converge to the global optima (episode return as ) while the vanilla MAPPO fails. However, note that CMAE is implemented on a tabular Q representation and suffers from exponentially increasing complexity, while our method can scale up well.
[…]
| Penalized Push-Box |
5.5 How Much Optimism Do We Need?
In this section, we perform an ablation study to examine how the performance change when we gradually change the degree of optimism, i.e. setting different values of in the Leaky ReLU extension. We experiment on the HalfCheetah 6x1 in MA-MuJoCo and Random 3 in Overcooked. Note that OptiMAPPO takes and degrades to MAPPO when . The results shown in Figure 4 indicate that optimism helps in both tasks to a wide range of degrees. While in the HalfCheetah 6x1 task, the performance improvement turns out clearly proportional to the degree of optimism. In Random 3 task, even when , OptiMAPPO converges as well as the full optimism. This can be partially due to the rewards in Overcooked being discretized in a coarse grain. Overall, the experiments augment our hypothesis that the optimism helps to overcome the relative overgeneralization in multi-agent learning and eventually helps to converge to a better joint policy.
5.6 Does Q-learning Based Optimism Work?
We empirically analyze whether Q-learning based optimism can boost performance in the Overcooked tasks using hysteretic Q-learning (Omidshafiei et al., 2017; Palmer et al., 2018). Figure 5 shows average Q values during training with different degrees of optimism. When decreases excessively, i.e. optimism increases too much, performance decreases. In both tasks shown in Figure 5, the highest optimism estimates the highest Q-values while showing the worst performance.
Different from the Q-learning based optimism methods, our proposed optimistic MAPG method performs the advantage estimation in an on-policy way and thus circumvents the overestimation problem naturally. Therefore, our method can fully employ the advantage of optimism and demonstrates strong performance in complex domains.
6 Limitation and Conclusion
Our optimistic updating approach yields state-of-the-art performance. However, as with other optimistic updating methods (Lauer & Riedmiller, 2000; Matignon et al., 2007), optimism can lead to a sub-optima when misleading stochastic rewards exist. The proposed Leaky ReLU extension allows further adaptive adjustment of the optimism degree to balance between optimism and neutrality and may allow to reduce the severity of stochastic rewards but this requires future investigation. In addition, lenient agents (Panait et al., 2006; Wei & Luke, 2016; Palmer et al., 2018) provide a set of heuristic techniques to adapt the degree of optimism, applicable to our method, to mitigate the problem with stochastic rewards. We leave this also as future work.
Motivated by solving the relative overgeneralization problem, we investigate the potential of optimistic updating in state-of-the-art MAPG methods. We first introduce a general advantage reshaping approach to incorporate optimism in policy updating, which is easy to implement based on existing MAPG methods. To understand the proposed advantage transformation, we provide a formal analysis from the operator view of policy gradient methods. The analysis shows that the proposed advantage shaping retains the optimality of the policy at a fixed point. Third, we extensively evaluate the instantiated optimistic updating policy gradient method, OptiMAPPO. Experiments on a wide variety of complex benchmarks show improved performance compared to state-of-the-art baselines with a clear margin.
Acknowledgements
We acknowledge the computational resources provided by the Aalto Science-IT project and CSC, Finnish IT Center for Science, and, funding by Research Council of Finland (353138, 327911, 357301). We also thank the ICML reviewers for the suggestions to connect our work to coordinated exploration methods.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Operator View of Policy Gradient
As shown in (Ghosh et al., 2020), the policy update in vanilla policy gradient can be seen as doing a gradient step to minimize
where is the discounted stationary distribution induced by the policy . and the distribution over actions are defined as
which corresponds to two successive operation by the projection operator and the improvement operator ,
The improvement operator tries to improve the policy into general function space via the information provided by the Q values, while the projection operator projects the into the policy function space . In this way, the policy gradient and Q-learning can be connected using the same polynomial operator , where the REINFORCE (Williams, 1992) is recovered by setting and Q-learning is obtained at the limit .
More sophisticated policy gradient methods arise by designing different which constructs different candidate distributions before being projected to the policy function space. For example, MPO (Abdolmaleki et al., 2018) uses a normalized exponential of Q values .
Appendix B Proof of Proposition 4.1
Proof of Proposition 4.1.
Following the proof for the regular policy gradient in (Ghosh et al., 2020), we replace the non-negative Q function with clipped advantages, which guarantees a valid probability distribution in the projection operators. Therefore, we have
where and are defined as:
Appendix C Implementation Details
We introduce the important implementation details here and the full details can be found in our code.
C.1 Pseudo Code for OptiMAPPO
C.2 Key Hyper-Parameters
Repeated Matrix Games:
In the two repeated matrix games, since the observation for each time step is fixed as a constant and the learned state value function is uninformative, we compute the advantage using TD(0) in both OptiMAPPO and MAPPO, instead of the GAE in order to reduce the noise.
MA-MuJoCo:
In all the tasks of MA-MuJoCo, we use the same hyperparameters listed in Table 5. The implementation is based on the HAPPO (Kuba et al., 2022) codebase, and the other hyperparameters are the default.
| Hyper-parameters | MAPPO | OptiMAPPO | HAPPO | HATRPO |
| Recurrent Policy | No | No | No | No |
| Parameter Sharing | No | No | No | No |
| Episode Length | ||||
| No. of Rollout Threads | ||||
| No. of Minibatch | ||||
| Policy Learning Rate | ||||
| Critic Learning Rate | ||||
| Negative Slope | N/A | N/A | N/A | |
| KL Threshold | N/A | N/A | N/A |
Overcooked:
In all the tasks of Overcooked, we use the same hyperparameters listed in Table 6.
Two example tasks in Overcooked are shown in Figure 6. In task Random3, the agents need to learn to circle through the corridor while avoiding blocking each other. In Unident_s, the agents learn to collaborate using their closest items to complete the recipe in an efficient way. However, each agent can also finish their own recipe without collaboration.
| Hyper-parameters | MAPPO | OptiMAPPO | HAPPO | HATRPO |
| Recurrent Policy | No | No | No | No |
| Parameter Sharing | Yes | Yes | No | No |
| Episode Length | ||||
| No. of Rollout Threads | ||||
| No. of Minibatch | ||||
| Policy Learning Rate | ||||
| Critic Learning Rate | ||||
| Negative Slope | N/A | N/A | N/A | |
| KL Threshold | N/A | N/A | N/A |
FACMAC:
We use the original code base from FACMAC (Peng et al., 2021) with default hyper-parameter.
Hysteretic DQN:
We implement Hysteretic DQN using the same state representation network as our method and select the best performance from three different values: on all the Overcooked tasks.
Appendix D More Results of Popular Deep MARL methods on Matrix Games
In order to show the RO problem in existing MARL methods, we cite the following results of common deep MARL methods on the penalty and climbing games from the benchmarking paper (Papoudakis et al., 2021). As shown in Table 7, popular MARL methods without explicitly considering the RO problem fail to solve the matrix games.
Appendix E Results of Optimistic MAA2C
We also apply the proposed advantage clipping to multi-agent advantage actor-critic (MAA2C) (Papoudakis et al., 2021), dubbed OptiMAA2C. Figure 7 demonstrates that OptiMAA2C improves MAA2C similarly to how OptiMAPPO improves MAPPO.
Appendix F Comparison with NA-MAPPO on MA-MuJoCo
NA-MAPPO (Hu et al., 2021) modifies the advantage estimates by injecting Gaussian noise in order to enhance the exploration of MAPPO. We compare our method with NA-MAPPO on two MA-MuJoCo tasks, HalfCheetah 6x1 and Ant 8x1. The result in Figure 8 shows that our method consistently outperforms NA-MAPPO.
Appendix G Full Results on MA-MuJoCo
We show the average and maximum return of evaluation episodes during training on all MA-MuJoCo tasks in Figure 9 and Figure 10, which shows our method outperforms the baselines on most tasks and matches the rest. With respect to the maximum episode return in Figure 10, our algorithm demonstrates clearer margins over the baselines.